Fermi National Laboratory
 

Accelerator Update for July 2, 2007 - July 4, 2007

 

Monday July 2

The day shift began with Operations monitoring the Tevatron (TeV) store 5531 (32.1E30) and the Recycle stash (162.5E10), with the Antiproton Source (Pbar) stacking (10.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI, MiniBooNE, and the Meson MTest area taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 9:27 AM, a MI RF station (MIRF11) tripped off and caused the NuMI beam permit to trip off. Everything reset.

At 11:14 AM, Operations reported that earlier the Switchyard (SY) beam permit tripped off due to bad beam positions at F36 caused by the QXR and would not reset. An expert suggested that operators adjust the gain for the QXR, which was what they did. MTest resumed taking beam.

Operations prepared the Linac at 1:29 PM for an access. Linac completed their access and resumed sending beam an hour later.

Operations reported at 3:55 PM that the Recycler had stashed antiprotons twice during the day shift.

The MTest experiment reported at 6:32 PM that they were done with the beam for the night.

Operations terminated store 5531 at 10:05 PM.

Operations put the TeV into shot setup at 10:17 PM.

At 10:42 PM, Operations reported that the MI-40 kickers tripped off. Operators reset them.

Operations reported at 11:59 PM that the Recycler had stashed three times during the evening shift.

Tuesday July 3

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (12.4E10), with Pbar stacking (51.7E10), with Recycler experts conducting studies, and with NuMI and MiniBooNE taking beam.

Operations established store 5532 at 12:21 AM with an initial luminosity of 197.72E30.

Operations halted all beam for a scheduled access to Booster and MI at 6 AM. They were both out of access by 9:53 AM.

Operations established beam to MiniBooNE at 11:50 AM.

Operations established beam to MTest at 1:14 PM, but then a SY problem halted their beam. A SY expert investigated. Operators adjusted the QXR gain and MTest resumed taking beam.

Pbar resumed stacking at 2:18 PM.

A MI expert began a slipstacking study at 2:43 PM.

Operations reported 3:58 PM that the Recycler was switching to its new damper system.

Operations reported at 3:58 PM that the Recycler had stashed once during the day shift.

MTest experimenters reported at 6:01 PM that they were done with beam for the night.

At 7:54 PM, the Pbar lithium lens tripped off on a ground fault indication. Operations contacted experts and then, per instruction, reset the lens power supply.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Wednesday July 4

The midnight shift began with Operations monitoring store 5532 (33.6E30) and the stash (206.6E30), with Pbar stacking (56.6E10), and with NuMI and MiniBooNE taking beam.

Operations terminated store 5532 at 2 AM.

Operations put the TeV into shot setup at 2:10 AM.

At 3:29 AM, Operations reported that they had been having problems with electron cooling for about 30 minutes. They turned off electron cooling for 10 minutes to resolve some MI beam position problems. After the cooling was turned back on and cooled for a while, Recycler beam resumed normal emittances.

Operations established store 5533 at 5:08 AM with an initial luminosity of 165.9E30.

Operations established beam to MTest at 6:02 AM.

Operations reported at 7:58 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/05/2007    email Fermilab